Graph a typical isoquant curve for the following production functions and determine whether they have convex indifference curves. Also calculate MRTS for each of the following functions
Q(x,y)=3x+y
Q(x,y)=√(x.y)
Q(x,y)=√x+ y
1
Expert's answer
2020-07-14T09:45:46-0400

a)


"Q(x,y)=3x+y"




"\frac {\delta Q}{\delta x}=3"

"\frac {\delta Q}{\delta y}=1"

"MRTS=\frac {3}{1}=3"

b)

"Q(x,y)=\sqrt {xy}"


"\frac {\delta Q}{\delta x}=\frac {y}{2 \sqrt {xy}}"

"\frac {\delta Q}{\delta y}=\frac {x}{2 \sqrt {xy}}"

"MRTS=\frac {y}{x}"

c)


"Q(x,y)=\sqrt {x+y}"


"\frac {\delta Q}{\delta x}=\frac {1}{2 \sqrt{x+y}}"

"\frac {\delta Q}{\delta y}=\frac {1}{2 \sqrt{x+y}}"


"MRTS=1"